Need help will pay.
Evaluate the expression.
P(7, 2) · P(4, 3)

Answer by ikleyn(52803)

.
Need help will pay.
Evaluate the expression.
P(7, 2) · P(4, 3)
~~~~~~~~~~~~~~~~~~~~~~~

If P(7,2), P(4,3) designates the number of permutations, then


P(7,2) = 7*6 = 42,


P(4,3) = 4*3*2 = 24,


and  P(7,2) · P(4,3)  = 42*24 = 1008.